OpenStudy (anonymous):

1/2(2x-3)+2/3(4x+1)=3-1/4(2x+2/3)

OpenStudy (mertsj):

Let's multiply each term fy 12 and thereby clear all the fractions.

OpenStudy (mertsj):

\[12\times\frac{1}{2}(2x-3)+12\times\frac{2}{3}(4x+1)=12\times3-12\times\frac{1}{4}(2x+\frac{2}{3})\]

OpenStudy (mertsj):

\[6(2x-3)+8(4x+1)=36-3(2x+\frac{2}{3})\]

OpenStudy (mertsj):

\[12x-18+32x+8=36-6x-6\]

OpenStudy (mertsj):

Perhaps you can take it from there.

OpenStudy (anonymous):

thank you so much

OpenStudy (mertsj):

yw